# KAN-powered large-target detection for automotive radar

### v1

@@ Abstract
This paper presents a novel radar signal detection pipeline focused on detecting large targets such as cars and SUVs. Traditional methods, such as Ordered-Statistic Constant False Alarm Rate (OS-CFAR), commonly used in automotive radar, are designed for point or isotropic target models. These may not adequately capture the Range-Doppler (RD) scattering patterns of larger targets, especially in high-resolution radar systems. Additional modules such as association and tracking are necessary to refine and consolidate the detections over multiple dwells. To address these limitations, we propose a detection technique based on the probability density function (pdf) of RD segments, leveraging the Kolmogorov-Arnold neural network (KAN) to learn the data and generate interpretable symbolic expressions for binary hypotheses. Beside the Monte-Carlo study showing better performance for the proposed KAN expression over OS-CFAR, it is shown to exhibit a probability of detection ($P_{\mathrm{D}}$) of $96\%$ when transfer learned with field data. The false alarm rate ($P_{\mathrm{FA}}$) is comparable with OS-CFAR designed with $P_{\mathrm{FA}}=10^{-6}$. Additionally, the study also examines impact of the number of pdf bins representing RD segment on performance of the KAN-based detection.
@@ I Introduction
Consequently, large automotive targets such as cars, SUVs, and trucks fall within this regime with characteristic lengths $>>10\lambda$. Traditionally, the automotive radar signal processing (RSP) pipeline assumes a point or isotropic target model, performing detection on a cell-by-cell basis. However, for large targets, this approach often results in multiple detections per acquisition, including both valid and spurious ones, influenced by the preset false alarm rate. Managing these detections in real time is challenging, thus requiring refinement and association, a task handled by the tracking module post-detection.
@@
The use of deep learning (DL), convolutional (CNN) and recurrent neural networks (RNN) have proven effective in representation learning by simplifying complex data [11]. Their application in radar systems for target detection is becoming more prevalent [12, 13, 14, 15]. In [12], CA-CFAR detections obtained with a reduced threshold (higher detections and false alarms) undergo target classification by a CNN, effectively maintaining increased detection rates while reducing false alarms. Truncated statistics are also augmented with neural networks (TS-NN CFAR) [16] to outperform OS-CFAR and traditional mean-level algorithms. LSTM in [17] utilizes the target scattering spread in the RD segment to classify and improve the detection of closely-spaced targets using CFAR detections as the midpoint. In the above techniques, neural networks are used along with CFAR, thus increasing the real-time computational load.
Studies have explored CNN-based binary detectors with sliding window sizes larger than the target echo in RD map [13]. A deep radar detector was developed to directly process the complex IF signal for target detection and localization [15]. PointNet-based CFAR detection has been applied to multi-object detection and sea clutter classification via segmentation of raw complex pulse echo data [18]. A VGGNet-inspired CNN, combined with region search and probability sorting using Intersection-over-Union (IoU), is used for target identification [19]. In  [14], Faster R-CNN is employed for object detection, along with particle filtering for enhanced tracking. However, these techniques exhibit higher complexity, and face challenges with segmentation labeling. Furthermore, the lack of explainability in neural networks raises regulatory concerns for automotive applications.
In this work, we approach the large-target detection problem with a Range-Doppler (RD) segment-based detection in contrast to the conventional cell-by-cell-based detection techniques. An RD segment detected to have a target has the merit of enriched scattering extent coverage, thereby enabling superior target representation.

### v2

@@ Abstract
This paper presents a novel radar signal detection pipeline focused on detecting large targets such as cars and SUVs. Traditional methods, such as Ordered-Statistic Constant False Alarm Rate (OS-CFAR), commonly used in automotive radar, are designed for point or isotropic target models. These may not adequately capture the Range-Doppler (RD) scattering patterns of larger targets, especially in high-resolution radar systems. Additional modules such as association and tracking are necessary to refine and consolidate the detections over multiple dwells. To address these limitations, we propose a detection technique based on the probability density function (pdf) of RD segments, leveraging the Kolmogorov-Arnold neural network (KAN) to learn the data and generate interpretable symbolic expressions for binary hypotheses. Beside the Monte-Carlo study showing better performance for the proposed KAN expression over OS-CFAR, it is shown to exhibit a probability of detection ($P_{\mathrm{D}}$) of $96\%$ when transfer learned with field data. The false alarm rate ($P_{\mathrm{FA}}$) is comparable with OS-CFAR designed with $P_{\mathrm{FA}}=10^{-6}$. The study also examines how the number of pdf bins in the RD segment affects the performance of KAN-based detection.
@@ I Introduction
Consequently, large automotive targets such as cars, SUVs, and trucks fall within this regime with characteristic lengths $>>10\lambda$. Traditionally, the automotive radar signal processing (RSP) pipeline assumes a point or isotropic target model, performing detection on a cell-by-cell basis. However, for large targets, this approach often results in multiple detections per acquisition, including both valid and spurious ones, influenced by the preset false alarm rate. Managing real-time detections is challenging, requiring refinement and association, which is handled by the tracking module after detection.
@@
The use of deep learning (DL), convolutional (CNN) and recurrent neural networks (RNN) have proven effective in representation learning by simplifying complex data [11]. Their application in radar systems for target detection is becoming more prevalent [12, 13, 14, 15, 16]. In [12], CA-CFAR detections obtained with a reduced threshold (higher detections and false alarms) undergo target classification by a CNN, effectively maintaining increased detection rates while reducing false alarms. Truncated statistics are also augmented with neural networks (TS-NN CFAR) [17] to outperform OS-CFAR and traditional mean-level algorithms. LSTM in [18] improves detection of closely-spaced targets with CFAR, while [19] uses lidar-based ground truth for extended targets with a Doppler encoder and 3D CNN. Both methods increase computational load and rely on external sensors or manual labeling, making large target detection challenging.
Studies have explored CNN-based binary detectors with sliding window sizes larger than the target echo in RD map [13]. A deep radar detector developed to process the complex IF signal for target detection and localization [15].PointNet-based CFAR detection has been applied to multi-object detection and sea clutter classification via segmentation of raw complex pulse echo data [20]. A VGGNet-inspired CNN, combined with region search and probability sorting using Intersection-over-Union (IoU), is used for target identification [21]. In  [14], Faster R-CNN is employed for object detection, along with particle filtering for enhanced tracking. However, these techniques exhibit higher complexity, and face challenges with segmentation labeling. Furthermore, the lack of explainability in neural networks raises regulatory concerns for automotive applications.
This work tackles large-target detection using Range-Doppler (RD) segment-based detection, unlike conventional cell-by-cell techniques. An RD segment with a detected target offers enriched scattering coverage for a better target representation.